OpenStudy (anonymous):

Find the equation of the plane that is parallel to the plane 5x-4y+z=1 and passes through the point (2,-1,-2)

OpenStudy (helder_edwin):

if u replace (x,y,z)=(2,-1,-2) then \[ \large 5x-4y+z=5\cdot2-4(-1)+(-2)=10+4-2=12 \] then the planes's equation is \[ \large 5x-4y+z=12 \]

OpenStudy (anonymous):

fundamental point to remeber here is PARALLEL PLANES HAVE SAME NORMALS So given a plane , find its normal, then use the point normal formula to find the equation of other plane

OpenStudy (anonymous):

does that make sense?

OpenStudy (anonymous):

Yes, psi9epsilon, thank you!